Tony had 13 bills in his wallet. They were all 5’s or 10’s. The total amount of money in his wallet was $90. How many 5’s in his wallet?

heres another ? thanxs soooooo much!!!!!!!!!!!!!!!!!
There were 200 tickets sold at the school play. Tickets were $2 for students and $3 for adults. The total amount collected was $490. How many tickets were sold?
thanx

Answer by stanbon(75887)

Tony had 13 bills in his wallet. They were all 5’s or 10’s. The total amount of money in his wallet was $90. How many 5’s in his wallet"
Let number of $5 bills be "x".
Then number of $10 bills must be "13-x".
Value of the $5 bills is 5x dollars
Value of the $10 bills is 10(13-x)=(130-10x) dollars
EQUATION:
value + value = $90
5x+130-10x=90
-5x=-40
x=8 (number of $5 bills)
13-x=5 (number of $10 bills)
-------------------------------------------------------
There were 200 tickets sold at the school play. Tickets were $2 for students and $3 for adults. The total amount collected was $490. How many tickets were sold?
Let number of $2 tickets be "x"
Then number of $3 tickets is "200-x"
Value of the $2 tickets is 2x dollars
Value of the $3 tickets is 3(200-x)= 600-3x dollars
EQUATION:
value + value = $490
2x + 600-3x=490
-x=-110
x=110 (number of $2 tickets sold)
200-x=90 (number of $3 tickets sold)
Cheers,
Stan H.